### v2

# DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection

Abstract
Infrared small target detection (ISTD) is widely used in civilian and military applications. However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve edge information of small targets. DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features. Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information. We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/DATransNet.
Index Terms:
Infrared small target detection (ISTD), convolution neural network (CNN), Dynamic Attention Transformer, global feature extraction.
I Introduction
INFRARED small target detection (ISTD) is critically important for numerous space-based computer vision applications. Although ISTD has been extensively studied, it remains a challenge. The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
Currently, the ISTD methods can be categorized into model-driven and data-driven methods. Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [1], enhance target visibility in complex backgrounds while suppressing background noise. 2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3], leverage human visual perception to improve local information. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and Rewight infrared patch tensor (RIPT) [5], decompose small targets and backgrounds using low-rank matrices. However, these techniques require manual parameter settings and often struggle to adapt to complex scenarios.
With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (ISTD). For example, the Asymmetric Context Modulation (ACM) network [6] introduces an asymmetric feature fusion technique, an alternative to conventional skip connections in U-Net. Building on this, DNANet [7] implements a multilayer nested architecture that supports progressive and adaptive interactions among feature layers. Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion. Additionally, GSTUnet [9] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets. Generally, these data-driven methods surpass traditional model-driven approaches.
Some challenges faced by data-driven methods, such as the substantial number of parameters and the potential loss of local feature information. Our motivation focus on enhancing the feature information from small targets.
Although Data-driven methods achieve good performance, they suffer from a substantial number of parameters and potential loss of local information. Therefore, we propose some methods to improve the ISTD network, and the new network is called Dynamic Attention Transformer Network (DATransNet). DATransNet incorporates two key innovation components to improve feature extraction. First, we propose the Dynamic Attention Transformer (DATrans), which utilizes Transformers to simulate the central difference convolution (CDC) and enable the network to retain the detailed information of the inputs in deep layers and give dynamic weights to differences result from differenct directions. Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information.
The main contributions of this letter are as follows.
  1. 1.
    We propose the DATransNet, which utilizes the DATrans to simulate the CDC with dynamic weights, extracting the edge and point information.
  2. 2.
    We propose the GFEM to improve the CNN from only focusing on local details to integrating these details with global information.
  3. 3.
    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms existing methods.
II Methods
The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
II-A Dynamic Attention Transformer (DATrans)
CDCs are defined in Eq. 1 and are employed to extract edge and gradient features. Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. $out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$. $w_{\hat{c}ijxy}$ is the weight applied to the difference $(o_{ixy}-b_{ijxy})$, which contributes to determining $out_{\hat{c}xy}$. A significant limitation of CDCs is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
$out_{\hat{c}xy}=\sum^{c_{inp}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
We can transform Eq. 1 into a fusion of Eq. 2 and Eq. 3. The process described in Eq. 2 is illustrated in Fig. 2. As shown in Fig. 2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{inp}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{inp}\times wh}$. In Eq. 3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ and the local difference matrix $D$ to obtain the final result $Out$. Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
$D=reshape(Conv_{3\times 3}(T))\$ (2)
$Out=M_{w}D$ (3)
Figure 2: $D$ is obtained by performing edge detection convolution on $T$ and unfolding the output.
When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness. Based on the analysis above, we introduce the DATrans to simulate CDC and extract the details using dynamic weights from low-level layers for ISTD. The structure of DATrans is shown in Fig. 3.
Figure 3: The various heads in DATrans employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
In the various heads of DATrans, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
$Out^{n}$ $=Softmax(Norm(\frac{Q^{n}(K^{n})^{T}}{w}))V^{n}$ (5)
$=M^{n}V^{n}=(M^{n}W_{V}^{n})D^{n}=M^{n}_{mix}D^{n}$
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$. According to Eq. 3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n. $Norm$ is the Instance Normalization. At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
$Out=Conv_{1\times 1}(reshape(Out^{1},...,Out^{m}))$ (6)
Overall, the gradient information from features at different scales can be fully extracted, enhancing the model’s ability to adapt to complex backgrounds.
II-B Global Feature Extraction Module (GFEM)
Background information is crucial alongside detailed features in ISTD. However, perceiving background relies on global features, which the DATrans struggles to handle effectively. To address this limitation and enhance global information perception, we propose the GFEM, as illustrated in Fig. 4.
In GFEM, we utilize the attention mechanism to facilitate the transmission of information across distant pixels. This module includes two key aspects to broaden perception: First, we employ the non-local attention mechanism [10] to compute the spatial attention, enabling connectivity between pixels over long distances. Second, we incorporate a squeeze-and-excitation block [11] to enhance channel-wise perception. Finally, we concatenate the outputs from the global spatial and channel attention modules and apply convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
Subsequently, each feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
Figure 4: The structure of the GFEM using the Non-local (Non-local Attention Module) to capture global spatial features and SE Block (Squeeze-and-Excitation Block) to extract global channel information.
II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of DATransNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq. 7.
$p_{i,j}=\frac{1}{1+e^{-v_{i,j}}}$ (7)
where $p_{i,j}$ denotes the Sigmoid output of $v_{i,j}$, the model’s prediction at pixel $(i,j)$. Moreover, the SoftIoU loss is defined as:
$loss=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
In this equation, $g_{i,j}$ denotes the ground truth value at pixel $(i,j)$, typically 0 or 1.
Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms. The Red bounding box are zoom in, facilitating to auxiliary our to evaluation detection results. Our DATransNet can achieve high probability detection rate, and achieve precision localization and segmentation of their shapes.
III Experiment and Analysis
This section first describes the experimental details and the evaluation metrics. Following with a series of ablation studies to verify the proposed modules. Finally, a comprehensive comparative experiments with other methods on qualitative and quantitative results demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
III-A Dataset and Evaluation Metrics
The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[12]. The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$). Meanwhile, we plot the ROC curves of various methods to evaluate the robust of DATransNet.
III-B Experimental Environment and Parameter Settings
Our model is implemented by PyTorch framework, then training on NVIDIA GeForce RTX 4080 GPUs. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
III-C Ablation Study
III-C1 Module-wise Performance Gain
In this ablation study, we begin with the most primitive Baseline, Which includes only feature extraction and upsampling with in the U-Net framework. Then, We test the performance of the DATransNet w/o each module, the results are shown in Tab. I. The results illustrate the effectiveness of each module. The performance of our network improves progressively with the inclusion of DATrans and GFEM. DATrans, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception. In ISTD, the capability to learn gradient information is more critical than access to global context.
TABLE I: Studies Of Different Components on NUDT-SIRST (The best results are bold)
Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
U-Net 91.31 95.44 97.98 4.46
U-Net+DATrans 94.25 97.03 98.83 2.73
U-Net+GFEM 92.32 96.14 96.30 3.96
U-Net+DATrans+GFEM 94.93 97.39 99.04 2.00
III-C2 Studies of Dilation Rate for DATrans
As mentioned in Section II-A, DATrans employs varying dilation rates across different detection heads. Tab. II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a uniform rate. Furthermore, networks with dilation ratios of 1 and 3 yield superior results compared to those with other dilation ratios. Therefore, the dilation ratio significantly impacts target detection, and we choose the ratios of 1 and 3 as the dilation ratios for our network.
TABLE II: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
Dilation Rate mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
1 93.53 96.30 98.89 5.58
1,2 94.41 97.12 98.83 2.46
1,3 94.93 97.39 99.04 2.00
1,5 94.24 97.03 98.65 1.47
1,2,3,4 94.58 97.20 98.04 2.21
TABLE III: Studies on GFEM on NUDT-SIRST (The best results are bold)
Non-local SE Block mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
$\checkmark$ -- 94.69 97.27 98.73 2.57
-- $\checkmark$ 94.53 97.19 98.23 1.75
$\checkmark$ $\checkmark$ 94.93 97.39 99.04 2.00
III-C3 Studies of GFEM
As mentioned in III, we conduct the experiments on the modules in the GFEM. As we can see, the network combining the two modules performs better.
III-D Comparsion with State-of-the-art (SOTA) Methods
To evaluate the effectiveness of our proposed model, we compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K datasets, including SCTransNet [16], UIUNet [8], ACM [6], ALCNet [14], AGPCNet [15], RDIAN [13], and DNANet [7]. As shown in Tab. IV, and our method performs excellently. We utilize DATransNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation. The comparison experiments include qualitative results and quantitative results, receptively.
III-D1 Qualitative Result
As shown in Fig. 5, our model preserves the shape of targets more closely to ground truth with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods, achieve satisfied results.
TABLE IV: Quantitative Comparsion With Different Methods on NUDT-SIRST and IRSTD-1K (The best results are bold, second best results are underline.)
Module Metrics NUDT-SIRST IRSTD-1K
Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
ACM[6] 0.40 70.97 82.99 97.67 7.33 63.78 77.91 93.93 5.81
RDIAN[13] 0.90 87.78 93.47 97.88 9.67 64.08 78.12 93.34 5.30
ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
AGPCNet[15] 13.36 85.38 92.11 97.88 5.95 67.54 80.64 94.27 4.76
DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96
UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 68.06 80.91 92.59 5.85
SCTransNet[16] 11.19 93.83 96.77 98.20 1.56 67.97 80.90 93.26 6.24
Ours 4.46 94.93 97.39 99.04 2.00 68.26 81.11 95.95 4.42
III-D2 Quantitative Results
As shown in the Tab. IV, our method achieves impressive performance, including an $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $2.00\times 10^{-6}$. Similarly, on the IRSTD-1K dataset, our method is a leading network.
Figure 6: ROC curves of varied networks on NUDT-SIRST datasets.
Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig. 6. Although UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak. In contrast, ALCNet performance shows a significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics. The ROC curve corroborates the findings in Tab. IV. The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
IV Conclusion
In this letter, we propose an ISTD network that enhances detection performance. The DATrans was proposed to enhance the local gradient feature extraction and detection performance of the network. Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing DATrans from overlooking background information due to its emphasis on local gradient differences. Then the local and global feature are fused. We conducted extensive experiments indicating satisfied results with fewer parameters. The model is a back-box approach because its capability for dynamic weight adjustments lacks mathematical proof. Moreover, it suffers from false alarms when processing high-noise images. To address these limitations, we aim to improve the model in two aspects. At first, we plan to improve the interpretability of the network. Then, we would like to integrate temporal information into the model to improve its robustness.